Parth (parthkohli):

Notation of a sequence: What do we mean by \(n \ge 1\) in \(\{ 2^n\}_{n \ge 1}\)? Is it all the terms of the sequence only have \(n = 1\) or more? So basically \(n\) iterates to the positive numbers.

Parth (parthkohli):

So, the sequence is \(\{2,4,8,16\cdots \}\).

Parth (parthkohli):

Am I correct?

Parth (parthkohli):

@agentx5 I need a genius for this.

OpenStudy (anonymous):

Powers of two, while n is greater than or equal to 1. Yep! You're good

OpenStudy (anonymous):

Infinite sequence in fact, right? (i.e.: what's the limit as n goes to infinity for 2\(^n\) ? ) ;-)

Parth (parthkohli):

Ah! Great stuff! \(\{ 1,2,3,4,5,6,7\} = \{n \} _{7 \ge n \ge 1}\)

Parth (parthkohli):

It never reaches anywhere, so it's infinity. That was an oral question :P

OpenStudy (anonymous):

This is how you typically see these though, as series (sums of the terms in a sequence) \(\sum_{n \ge 1}^{\infty} 2^n\) = \(\infty\) or "Diverges" as they say

Parth (parthkohli):